OpenStudy (anonymous):

Simplify: (5x2 + 3x + 4) + (2x2 - 6x + 3) 7x2 + 9x + 7 7x2 + 3x + 7 7x2 - 3x + 1 7x2 - 3x + 7

OpenStudy (anonymous):

medal

OpenStudy (anonymous):

@Loser66

OpenStudy (anonymous):

So you need to combine like terms to solve this. They are 5x^2+2x^2, 3x-6x, and 4+3. Do you know how to combine those?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

To do that, remove the letters, and do 5+2=7, 3-6=-3 4+3=7 Now put the letters back in, and you have 7x^2-3x+7 Do you understand it?

OpenStudy (anonymous):

yes i do now thank you so much @zeinas.

OpenStudy (anonymous):

You're welcome